A 17 mL solution of KOH has a concentration of 0.672 M. Calculate the pH and then
calculate the mass in grams of KOH in solution
1
Expert's answer
2020-05-15T10:23:20-0400

1)   pH = - log [H3O+]

pH = - log (0.672) = -(-0.17)=0.17

2)   m=M*Mr=0.672*56=37,6

37,6 g KOH in 100g of water

X g KOH in 17g of water

X= 37.6*17/100=6.3g